Hassock Corp. produces woven wall hangings. It takes 2 hours of direct labor to produce a single wall hanging. Hassock’s standard labor cost is $17 per hour. During August, Hassock produced 14,000 units and used 28,160 hours of direct labor at a total cost of $476,020. What is Hassock’s labor rate variance for August?

Multiple Choice

  • $2,685 favorable.

  • $2,700 unfavorable.

  • $2,700 favorable.

  • $5,385 favorable.

  • $2,685 unfavorable.
Answer #1

Labor rate variance

= (SR - AR) *AH

= (17 - 476,020/28,160) * 28,160

= (17 - 16.90) * 28,160

= 2700 Favourable